OpenStudy (anonymous):

Solve x2 + 18 = –9x.

OpenStudy (anonymous):

you first step would be to add 9x to both sides, giving you: x^2+9x+18=0 Now you next step is to think of two numbers that multiply to be 18 and add to 9, that way you can factor this quadratic

OpenStudy (lilg132):

x^2 + 9x + 18 = 0 (x + 6) (x + 3) = 0 x = -6 x = -3 remember to click good answer

OpenStudy (anonymous):

wow, lilg132 you can solve this types of problems, good job

OpenStudy (lilg132):

i thought you needed help lagrange, your welcome :p